### v1

# Towards Competitive Search Relevance For Inference-Free Learned Sparse Retrievers

Abstract.
Learned sparse retrieval, which can efficiently perform retrieval through mature inverted-index engines, has garnered growing attention in recent years. Particularly, the inference-free sparse retrievers are attractive as they eliminate online model inference in the retrieval phase thereby avoids huge computational cost, offering reasonable throughput and latency. However, even the state-of-the-art (SOTA) inference-free sparse models lag far behind in terms of search relevance when compared to both sparse and dense siamese models. Towards competitive search relevance for inference-free sparse retrievers, we argue that they deserve dedicated training methods other than using same ones with siamese encoders. In this paper, we propose two different approaches for performance improvement. First, we introduce the IDF-aware FLOPS loss, which introduces Inverted Document Frequency (IDF) to the sparsification of representations. We find that it mitigates the negative impact of the FLOPS regularization on search relevance, allowing the model to achieve a better balance between accuracy and efficiency. Moreover, we propose a heterogeneous ensemble knowledge distillation framework that combines siamese dense and sparse retrievers to generate supervisory signals during the pre-training phase. The ensemble framework of dense and sparse retriever capitalizes on their strengths respectively, providing a strong upper bound for knowledge distillation. To concur the diverse feedback from heterogeneous supervisors, we normalize and then aggregate the outputs of the teacher models to eliminate score scale differences. On the BEIR benchmark, our model outperforms existing SOTA inference-free sparse model by 3.3 NDCG@10 score. It exhibits search relevance comparable to siamese sparse retrievers and client-side latency only 1.1x that of BM25.
Passage retrieval, learned sparse retriever, knowledge distillation
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Retrieval models and rankingccs: Computing methodologies Natural language processing
1. Introduction
Information retrieval(IR) and question answering(QA) are fundamental tasks in the realm of information processing, widely employed in various web applications. Lexical-based algorithms such as TF-IDF and BM25 were once the dominant approach. These algorithms utilize inverted indexes, which was proven to be efficient. However, due to issues such as vocabulary mismatch (Zhao and Callan, 2010) and the lack of contextual information, their semantic retrieval capabilities are limited. In contrast, siamese dense retrievers have overcome the limitations of traditional lexical-based methods and have become the mainstream approach for semantic retrieval (Reimers and Gurevych, 2019). Nonetheless, the ANN algorithm requires a substantial amount of memory leading to a significant trade-off between search relevance and resource consumption (Malkov and Yashunin, 2018; Jegou et al., 2010). The interpretability of dense retrievers is also questioned. Recent years, learned sparse retrieval is proposed to address this obstacle and gained increasing attention (Dai and Callan, 2020; Formal et al., 2021b, a). This approach predicts token weights based on their semantics with context information. It expands token set with generative models (Nogueira et al., [n. d.], 2019) or masked language model heads (Bai et al., 2020; Zhao et al., 2021; Formal et al., 2021b, a, 2022; Lassance and Clinchant, 2022; MacAvaney et al., 2020; Lassance et al., 2024), thereby addressing the vocabulary mismatch problem. Since sparse embeddings can be integrated with inverted indexes, the retrieval process of learned sparse models is highly efficient without compromising recall. Moreover, learned sparse retrieval offers better interpretability because the contribution of each token can be intuitively understood by human.
Among the learned sparse models, the inference-free architecture is particularly attractive to search applications. This architecture degenerates online model inference for query encoding into simple tokenization, significantly reducing end-to-end search latency and the associated model deployment costs.
Early work like DEEPCT (Dai and Callan, 2020) and Doc2Query (Nogueira et al., [n. d.]) attempted to associate additional information with the original documents. However, search relevance could not be trained in an end-to-end manner. Proposed by Formal et al. (2021a), the SPLADE-doc architecture has achieved SOTA performance among inference-free retrievers. It predicts the token weights and expands the tokens with similar semantics. The search relevance and sparsity are tuned via end-to-end training. However, even the latest SOTA inference-free model, SPLADE-v3-Doc (Lassance et al., 2024), exhibits a significant gap in search relevance when compared with siamese sparse retrievers. On the BEIR benchmark, the average NDCG@10 score of SPLADE-v3-Doc is 4.7 lower than siamese sparse retrievers of the same size and training method. This disparity hinders its application in actual production environments.
In this paper, we focus on improving the search relevance of the inference-free sparse retriever through more effective training methodologies. The first challenge lies in the uniform penalty applied to all tokens by the FLOPS regularization (Paria et al., 2020).
The purpose of the FLOPS loss is to eliminate unnecessary tokens to sparsify the document representation. However, the uniform penalty from FLOPS loss has a side effect: it prevents the model from assigning tokens with large weights. Consequently, for semantically rich tokens that are important to be retained in the representation, their importance will be underestimated. To mitigate this effect, we propose a simple yet effective enhancement called IDF-aware FLOPS. This loss adjusts the penalty imposed by the FLOPS loss on different tokens, making the penalty inversely proportional to the token’s IDF value. It reduces the regularization penalty on unique tokens and increases the penalty on frequent tokens.
Through experiments, we demonstrate that IDF-aware FLOPS effectively improves the search relevance of the inference-free sparse retriever, and we discover that it effectively reduces the average FLOPS number for the retrieval process.
Subsequently, the pre-training phase is also explored in this paper. We argue that although the commonly used contrastive InfoNCE loss (Chen et al., 2020) is able to enhance the alignment and uniformity of representations (Wang and Isola, 2020), these two targets are not applicable in inference-free models. Because for inference-free models, all semantics are only encoded at the model-side. And the search relevance can not be improved by aligning the document representation with the bag-of-word query representation. In contrast, knowledge distillation presents a more suitable approach for the training (Formal et al., 2021a). Hofstätter et al. (2020) proposed to train dense retrievers by conducting knowledge distillation from cross-encoder rerankers, and Formal et al. (2022, 2021a) applied this method to the fine-tuning of sparse retrievers. However, for large-scale pre-training datasets, the inference workload of the teacher model can reach 10 times or even larger. And the inference cost of cross-encoders is impractical at these settings, especially when in-batch negatives are utilized. In this paper, we propose to build a strong teacher model by assembling siamese dense retrievers and siamese sparse retrievers. Siamese retrievers have a heterogeneous and superior architecture compared with the inference-free architecture. And their inference cost is applicable for large-scale pre-training. Moreover, the ensemble of dense and sparse retrievers can further enhance the upper bound of knowledge distillation, enlarging the space for performance improvement of our model. During the assembling process, we normalize the scores for heterogeneous retrievers. This prevents one retriever from dominating the assembled result, further balancing the contribution of teacher models.
We conduct experiments on 13 public datasets from the BEIR benchmark, and our model outperforms the existing SOTA inference-free sparse model by 3.3 average NDCG@10 scores. Its performance even surpasses many strong siamese retrievers. Our contributions can be summarized as follows: (1) We propose IDF-aware FLOPS, which effectively improves the search relevance and efficiency of inference-free sparse models. (2) We explore how to effectively pre-train inference-free sparse models and propose the ensemble teacher model of heterogeneous siamese models, which has reasonable inference costs and strong performance. (3) The zero-shot performance of our model outperforms the SOTA inference-free retriever by 3.3 NDCG@10 score. It also surpasses strong siamese retrievers including SPLADE-v3-DistilBERT and ColBERTv2. While its client-side latency is only 1.1x that of BM25.
2. Related Work
2.1. Dense Retrieval
Recent years, the use of language models to generate dense embeddings for text representations has become prevalent in QA and IR (Reimers and Gurevych, 2019; Conneau et al., 2017; Karpukhin et al., 2020; Qu et al., 2021). Continuous efforts have been made to improve the training methodologies for dense retrieval models, such as negative sampling (Qu et al., 2021; Zhan et al., 2021) and knowledge distillation (Hofstätter et al., 2020, 2021; Lin et al., 2021). To enhance the generalization capability of dense retrieval models, numerous studies explore pre-training techniques on text embeddings. Some works (Gao and Callan, 2021, 2022; Xiao et al., 2022) design auxiliary tasks to enrich the dense embeddings from models, while another tributary of previous work pre-train the model directly on constructed text pairs (Li et al., 2023; Chen et al., 2024; Wang et al., 2022), including unsupervised and weakly supervised data.
Knowledge distillation (Gou et al., 2021) utilizes the soft labels from teacher models to facilitate a more effective training process for student models so as to improve the accuracy. Researchers strive to select teacher models with inherent advantages, such as larger parameters size or superior architectures, to ensure the best possible knowledge transfer. Hofstätter et al. (2020) proposes using a cross-encoder reranker as the teacher model for the siamese dense retriever during fine-tuning. However, in the context of pre-training, the significantly larger data volume makes the use of cross-encoders prohibitively expensive.
Pre-training is a widely adopted technique to enhance the accuracy and generalization capabilities of dense retrievers. The contrastive InfoNCE loss is applied to massive amounts of unsupervised or weakly-supervised data. Previous work (Wang and Isola, 2020) illustrates that the contrastive InfoNCE loss improves the dense representation from the perspectives of alignment and uniformity. However, for the inference-free architecture, the asymmetric sparse representation is unaware of the query distribution, and the token weight is simply an importance measure for that token. Consistent with our experiments, pre-training with the InfoNCE loss does not improve the model as expected on dense retrievers.
The challenges associated with knowledge distillation and pre-training methodologies limit the application of these techniques to inference-free sparse retrievers.
2.2. Sparse Retrieval
Learned sparse retrievers have gained increasing attention due to their ability to perform semantic search while retaining the advantages of traditional lexical-based retrieval methods. DEEPCT (Dai and Callan, 2020) employs the BERT model to fit the token weights computed by heuristic rules. It modifies the term frequency, thereby influencing the match score in the BM25 algorithm. However, DEEPCT does not address the vocabulary mismatch issue inherent in lexical-based retrieval. Doc2query (Nogueira et al., 2019) and docTTTTTquery (Nogueira et al., [n. d.]) tackle this issue by using generative models to predict potential queries for the original document. These queries are indexed together with the original document, enabling the matching of tokens not present in the document. These methods represent a preliminary exploration of inference-free learned sparse retrievers. Nevertheless, these models are unable to apply ranking loss in an end-to-end manner, thereby limiting their performance.
Following these research studies, more works have been proposed to train sparse retrievers in an end-to-end manner. Ranking losses such as infoNCE which are already widely used in training dense retrievers are discovered to be also eligible for sparse retrievers. Methods such as SparTerm (Bai et al., 2020), EPIC (MacAvaney et al., 2020), and SPARTA (Zhao et al., 2021) predict token weights based on the estimation of token importance and then apply various sparsification strategies, such as top-k pooling or learned gating, to obtain sparse embeddings. The match score between the query and document is obtained by taking the inner product of their sparse embeddings. Since the output space is identical to the vocabulary space, tokens that do not appear in the document can also be expanded in the representation. However, these methods are not of an effective end-to-end optimization manner and did not combine ranking loss and sparsity simultaneously.
The SPLADE-series models (Formal et al., 2021b, a, 2022; Lassance et al., 2024) overcome this obstacle by introducing the FLOPS regularizer to the loss function, enabling an end-to-end integration of ranking loss and sparsity. Inspired by Hofstätter et al. (2020), they employ hard negatives and knowledge distillation from cross-encoders to improve model performance. Among the SPLADE-series models, SPLADE-doc (Formal et al., 2021a) performs model inference solely on documents and sums up the weights for all matched tokens as the match score. It eliminates the need for model inference during retrieval, and the model training is completely end-to-end. However, even SPLADE-doc offers SOTA performance among inference-free models, there is still a significant gap persists between it and siamese dense/sparse retrievers in terms of search relevance.
3. Preliminary
Our work is built upon the SPLADE-doc-distill [^1] The original paper (Formal et al., 2021a) proposes the SPLADE-doc architecture and knowledge distillation training method. We incorporate the knowledge distill on SPLADE-doc and call it SPLADE-doc-distill. We implement this baseline in the experiment session. model. In this section, we’ll introduce the details of the baseline method.
3.1. Model
SPLADE-doc-distill is an expansion model that predicts the estimated importance of tokens in the vocabulary space. The model reuses the masked language model head to generate a similarity distribution over the token vocabulary. Utilizing BERT as its backbone, the model operates with a sparse vector of 30,522 dimensions. It utilizes max pooling to aggregate the predictions across different positions and uses ReLU activation to sparsify the representation and ensure the weights are non-negative. SPLADE-doc-distill has an asymmetric architecture, which means it only performs model inference on documents. For an input document, let $w_{i,j}$ represent the weight of token $j$ at position $i$. The formula to get the sparse representation is as follows:
(1) $w_{j}=\max_{i}\log(1+\text{ReLU}(w_{i,j}))$
The sparsity arises from the FLOPs regularizer during training and the ReLU activation. For query representation, it can be illustrated as binary Bag-of-Words. That is, the weight $w_{j}$ of token $j$ in query representation is set to 1 only if the token $j$ is present in the query; otherwise, $w_{j}$ is set to 0. Consequently, the ranking score for a given query and document can be computed as the inner product of their sparse representations, which is formulated as follows:
(2) $s(q,d)=\sum_{j\in q}w_{j}^{d}$
3.2. Training
The optimization objective of SPLADE combines the ranking loss $\mathcal{L}_{\text{rank}}$ and sparse regularization loss $\mathcal{L}^{d}_{\texttt{FLOPS}}$, ensuring an end-to-end training approach that simultaneously optimizes the search relevance and representation sparsification. For SPLADE-doc-distill, the training loss $\mathcal{L}$ is formulated as follows:
(3) $\mathcal{L}=\mathcal{L}_{\texttt{rank}}+\lambda_{d}\mathcal{L}^{d}_{\texttt{FLOPS}}$
(4) $\mathcal{L}^{d}_{\texttt{FLOPS}}=\sum_{j\in V}{\bar{a}}_{j}^{2}=\sum_{j\in V}\left(\frac{1}{N}\sum_{i=1}^{N}w_{j}^{(d_{i})}\right)^{2}$
where $\lambda_{d}$ is the weight for FLOPS regularizer and $\overline{a}_{j}$ is the average weight of token $j$ in the batch. The FLOPS regularizer penalizes the dimensions with high average weights in the representation to achieve the sparsity. In the context of text retrieval, tokens occurs more frequently and tokens with larger weights get more penalty. For given triplets $(q,d^{+},d^{-})$ consisting of a query, a positive document, and a hard negative document, SPLADE-doc-distill utilizes the KL divergence function to fit the match score differences produced by the model on the given triplets to the predicted results of the cross-encoder teacher model, thereby performing knowledge distillation.
4. Method
4.1. IDF-aware FLOPS
For the vanilla SPLADE-doc-distill model, according to Equation 4, we note that the FLOPS loss imposes a uniform penalty on different tokens. This penalty increases in a quadratic way with the average weights of token $j$, while the growth rate of the match score on token $j$ is linear. Although the initial intention of the FLOPS regularizer was to sparsify the representation, one unintended consequence is that it prevents the model from generating excessively large weight for a single token. For semantically important tokens that are always retained once they occurs, FLOPS loss put an extra restriction on their weights.
Therefore, we propose the IDF-aware FLOPS, which is a simple yet highly effective approach. We multiply the model output with token IDF values as the token weight. When computing the FLOPS loss, we exclude the IDF values and only apply constraints on the model output. The IDF values are involved in the calculation of the match score during both training and inference phases. Formally, we modify the token weight and FLOPS loss formulas for Equation 1 and Equation 4 to the following:
(5) $w_{j}=\text{IDF}_{j}\cdot\max_{i}\log(1+\text{ReLU}(w_{i,j}))$
(6) $\mathcal{L}^{d}_{\texttt{FLOPS}}=\sum_{j\in V}{\bar{a}}_{j}^{2}=\sum_{j\in V}\left(\frac{1}{N}\sum_{i=1}^{N}\frac{w_{j}^{(d_{i})}}{\text{IDF}_{j}}\right)^{2}$
where the other formulas and procedures remain unchanged.
This approach ensures that the penalty imposed by the FLOPS regularization on the model outputs is inversely proportional to the token’s IDF value. This compensates for the potential underestimation of important tokens. Furthermore, IDF-aware FLOPS drastically increases the penalty on low-IDF tokens, efficiently reducing the complexity of the retrieval process.
To present a straightforward view on the typical low-IDF and high-IDF tokens, we conduct a case study on the MS MARCO dataset [^2] https://github.com/microsoft/MSMARCO-Passage-Ranking. We utilize the uncased BERT tokenizer, and the outcome is presented in Table 1. According to the results, tokens with the highest IDF values represent unique semantics. We should always keep them by giving less penalty. In contrast, tokens with the lowest IDF values are more akin to stopwords. Imposing a uniform regularization penalty on these tokens would be an unreasonable approach.
Table 1. Case study of IDF values and tokens on the MS MARCO dataset. We only consider tokens appearing in more than 1000 documents.
Lowest IDF Token IDF Value Highest IDF Token IDF Value
. 0.016 janeiro 9.086
the 0.135 yemen 9.086
, 0.176 interiors 9.086
of 0.298 underwood 9.086
and 0.318 trevor 9.085
As a widely subscribed factor in relevance calculation (such as TF-IDF or BM25), the IDF values are tracked real-time in most of the lexical search engines. And this provides the convenience of obtaining the IDF values. Furthermore, We find that precomputed IDF values derived from the training set also exhibit strong generalization capabilities. Details can be found in Section 5.4.1.
The standard SPLADE-doc model assigns a weight of 1 to all query tokens. By decoupling the IDF values and the model outputs, we can utilize the IDF values as query token weights when constructing the query body in the search engine. This enables the application of heuristic optimization rules, further enhancing search latency and throughput. Details can be found in Section 5.3.2.
Table 2. The list of datasets used in pre-training.
Dataset Type Number of Training Tuples Remaining Queries After Filtering
S2ORC (Lo et al., 2020) (Title, Abstract) pairs 41,769,185 500,000
WikiAnswers (Fader et al., 2014) duplicate questions 77,427,422 1,000,000
GOOAQ (Khashabi et al., 2021) (Question, Answer) pairs 3,012,496 2,274,901
SearchQA Question + Top5 text snippets 117,220 116,933
Eli5 (Fan et al., 2019) (Question, Answer) pairs 325,475 168,652
WikiHow (Koupaee and Wang, 2018) (Summary, Text) pairs 128,542 111,987
SQuAD (Rajpurkar et al., 2018) (Question, Answer_Passage) pairs 87,599 84,505
Stack Exchange (Title, Title) pairs of duplicate questions 304,525 132,490
(Body, Body) pairs of duplicate questions 250,519 98,266
(Title+Body, Title+Body) pairs of duplicate questions 250,460 117,839
Yahoo Answers (Zhang et al., 2015) (Title, Answer) pairs 1,198,260 361,312
(Question, Answer) pairs 681,164 139,584
(Title, Question) pairs 659,896 252,823
4.2. Ensemble Heterogeneous Knowledge Distillation
Pre-training is a widely adopted technique to improve the performance and generalization of retrieval models. To further boost the search relevance, we pre-train the model on an extensive corpus encompassing both unsupervised and weakly-supervised datasets. Subsequently, we fine-tune the model on a high-quality labeled dataset, specifically the MS MARCO dataset. To construct an effective optimization objective, we employ knowledge distillation techniques. The primary challenge lies in generating efficient supervisory signals for the large-scale pre-training data. For an input batch containing $N$ queires and $M$ documents, the cross-encoders need to inference at $O(MN)$ complexity. And the cost becomes impractical as data scales.
In this paper, we introduce a novel technique that leverages an ensemble of heterogeneous models as the teacher model for knowledge distillation on large-scale data, the assembling procedure is illustrated in Figure 1. Siamese dense and sparse retrievers are combined to generate supervisory signals for the input data, and we employ the KL Divergence loss function to transfer knowledge to the student model.
Figure 1. The procedure for pre-training with ensemble heterogeneous knowledge distillation.
Heterogeneous teacher models. The cross-encoders are used by previous work (Formal et al., 2022) to provide supervision signals. However, their inference cost is impractical to be applied on large-scale pre-training. In contrast to inference-free retrievers, siamese retrievers possess a superior architecture and significantly lower inference costs compared with cross-encoders. To compensate the accuracy drop compared to cross-encoders, we propose an ensemble approach that combines dense and sparse retrievers as the teacher model. Sparse and dense teachers emit different document recall process, forming a heterogeneous distillation framework. Research studies (Chen et al., 2024; Bühlmann, 2012) and industrial practices [^3] https://opensearch.org/blog/semantic-science-benchmarks/ demonstrate that combining these predictors results in a significantly more robust retriever. Consequently, the teacher model comprising an ensemble of dense and sparse retrievers is both efficient and accurate simultaneously. But there are still challenges for combining dense and sparse retrievers, as their match scores have disparate scaling. If not addressed properly, the retriever with a larger score scale may dominate the combined results, leading to biased or skewed outputs. For each retriever the scores are normalized before integration. We employ min-max normalization to scale all scores to the range of [0,1]. Subsequently, we combine the two retrievers by calculating their arithmetic mean and get the weighted sum. We then multiply the combined scores with a constant $S$ to scale them back for knowledge distillation. These steps ensure a balanced contribution from both sides, which are crucial for combining diverse retrieval models. The aforementioned process can be represented by the following formula:
(7) $\hat{s_{i}^{j}}=\frac{s_{i}^{j}-min(s^{j})}{max(s^{j})-min(s^{j})}$
(8) $\hat{s}=S\cdot\sum_{j}w^{j}s^{j}$
where $s_{i}^{j}$ are the match scores from model $j$ on document $i$, $w_{j}$ is the weight for model $j$, and $\hat{s}$ are the final supervisory signals. In our experiments, we use equal weight for the dense and sparse teacher model.
We use the KL Divergence to compute the loss for the ensemble scores and the output of our inference-free sparse model. The FLOPS regularizer is also applied during the pre-training phase. We utilize a small coefficient $\lambda_{d}$ for the FLOPS regularizer. The first reason is that for the inference-free sparse model, the involved tokens are limited, and we aim to avoid omitting any token during the pre-training phase. Secondly, we need to search the optimal hyperparameter $\lambda_{d}$ to strike a balance between search relevance and retrieval cost through multiple experiments. It will be more efficient to conduct these experiments solely for fine-tuning.
Data preparation. Regarding the pre-training phase, we use a subset of training data collected by Sentence Transformers.  [^4] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2#training-data The pre-training datasets were presented in the form of pairs, such as (Question, Answer), (Title, Content), and duplicate question/answer sets collected by the content providers. They are constructed either through automatic rules or human annotation, covering multiple domains. To harness the full potential of knowledge distillation, self-mined hard samples are used during training. We first train an inference-free sparse retriever without the pre-training phase as the miner model. Subsequently, for each query we use this model to mine the top $M$ relevant documents from the full documents collection of the source dataset. In each pre-training step, we commence by randomly selecting a dataset, followed by the random sampling of $N$ queries and their corresponding hard samples. Inspired by Wang et al. (2022), we use a consistency-based filtering approach to retain only those training samples where the labeled positive document is ranked among the top-$k$ retrieved documents.
Subsequent to the pre-training phase, we fine-tune the model on the MS MARCO dataset. Following Formal et al. (2021a), cross-encoders are utilized as teacher models during fine-tuning. We ensemble the cross-encoders with dense and sparse teacher models used in pre-training. At fine-tuning stage, we accomplish the final sparsification of the representation.
5. Experiments
5.1. Settings
5.1.1. Training Data
For the pre-training phase, we utilized a subset of the datasets collected by the Sentence Transformers project. Detailed datasets are listed in Table 2. Following the last paragraph in Section 4.2, samples where the positive document is not ranked among the top 10 results are filtered out. In each training step, we take the positive document and 7 hard negative documents for every query. For the S2ORC and WikiAnswers datasets, we only sample a portion of them to prevent the large sample size from dominating the pre-training dataset. Ultimately, there are 5359292 queries and their corresponding hard negative documents. For the fine-tuning phase, we utilize the MS MARCO passage ranking dataset. This dataset comprises 8,841,823 passages and 502,939 queries in the training set. For each query, we sample the top 100 hard negative documents to facilitate knowledge distillation.
5.1.2. Model Training
Co-Condenser (Gao and Callan, 2022) [^5] https://huggingface.co/Luyu/co-condenser-marco is hired as the backbone, which is of the same size with BERT-base model. The IDF values for tokens are calculated using the documents of MS MARCO dataset. If a token is not present in the dataset, its value is set to 1. The query IDF representation remains frozen throughout the training and evaluation processes.
For the teacher models of the pre-training phase, we employ SOTA dense and sparse retrievers , namely gte-large-en-v1.5666https://huggingface.co/Alibaba-NLP/gte-large-en-v1.5 and opensearch-neural-sparse-encoding-v1777https://huggingface.co/opensearch-project/opensearch-neural-sparse-encoding-v1. The total number of pre-training steps is 150,000. For each step, we sample 48 queries and 8 hard samples(1 positive document and 7 hard negative documents) for every query. We set the learning rate to 5e-5 and the FLOPS $\lambda_{d}$ to 1e-7. The max input length is set to 128. The constant $S$ is set to 10 for scaling back the assembled scores. In the fine-tuning phase, we utilize an ensemble teacher model comprising the two siamese retrievers used in the pre-training phase, as well as two cross-encoder re-rankers [^8] We’re using https://huggingface.co/castorini/monot5-3b-msmarco-10k and https://huggingface.co/cross-encoder/ms-marco-MiniLM-L-12-v2 . The total number of training steps is 50,000. For each step, we sample 40 queries and 10 hard negatives for each query. We set the learning rate to 2e-5 and the FLOPS $\lambda_{d}$ to 0.02. The max input length is set to 256. And the IDF values are calculated using the documents of MS MARCO dataset. The constant $S$ is set to 30 for scaling back the assembled scores.
5.1.3. Indexing and Evaluation
In this paper, a lexical search engine called OpenSearch [^9] https://opensearch.org/, is employed to construct the inverted index and perform the retrieval process as well. By leveraging the OpenSearch neural sparse feature, we can seamlessly integrate the writing and searching processes for custom learned sparse models. For evaluation metrics, we use the implementation of BEIR python toolkit to caculate the MRR, NDCG and recall rate. During evaluation, we use the IDF values derived from the MS MARCO dataset. The max input length is set to 512.
5.2. Search Relevance Evaluation
5.2.1. In-domain Performance
As we fine-tune the model on MS MARCO dataset, we report the in-domain performance on this dataset. Following Formal et al. (2021b), we report the MRR@10 and Recall@1000 on MS MARCO dev set. We also report the NDCG@10 and Recall@1000 for the TREC DL 2019 evaluation set [^10] TREC DL 2019 contains 43 queries for MS MARCO corpus annotated by human.. We train the SPLADE-doc model using knowledge distillation techniques described by Formal et al. (2021a) as baseline, and reference it as SPLADE-doc-distill in other sections. The results for baseline models are extracted from corresponding papers. The baseline models contain siamese dense encoders including ANCE (Xiong et al., 2020), TCT-ColBERT (Lin et al., 2021), ColBERTv2 (Santhanam et al., 2022), RocketQA (Qu et al., 2021), RocketQAv2 (Ren et al., 2021), CoCondenser (Gao and Callan, 2022), TAS-B (Hofstätter et al., 2021). And we also include sparse retrievers including BM25, SparTerm (Bai et al., 2020), DEEPCT (Dai and Callan, 2020), doc2query-T5 (Nogueira et al., [n. d.]) and SPLADE-series models (Formal et al., 2021b, a, 2022; Lassance et al., 2024). SPLADE-v3-Doc applies knowledge distillation on SPLADE-doc. It also employs tricks to improve the model’s in-domain search relevance, and these tricks are orthogonal to what we proposed in this paper. The results are shown in Table 3.
Table 3. Evaluation result on MS MARCO dataset. Models marked with † means the model is trained and evaluated by us.
Model MS MARCO dev TREC DL 2019
M@10 R@1000 NDCG R@1000
Dense Retrievers
ANCE 33.0 95.9 64.8 -
TCT-ColBERT 35.9 97.0 71.9 76.0
ColBERTv2 39.7 98.4 - -
RocketQA 37.0 97.9 - -
RocketQAv2 38.8 98.1 - -
CoCondenser 38.2 98.4 - -
TAS-B 34.7 97.8 71.7 84.3
Sparse Retrievers
SparTerm 27.9 92.5 - -
DistilSPLADE-max 36.8 97.9 72.9 86.5
SPLADE-v3-DistilBERT 38.7 - 75.2 -
Inference-free Sparse Retrievers
BM25 18.4 85.3 50.6 74.5
DeepCT 24.3 91.3 55.1 75.6
doc2query-T5 27.7 94.7 64.2 82.7
SPLADE-doc 32.2 94.6 66.7 94.7
SPLADE-doc-distill† 36.5 96.9 69.8 74.2
SPLADE-v3-Doc 37.8 - 71.5 -
Our Model† 37.8 97.5 72.1 79.8
Table 4. Evaluation result on BEIR dataset. Models marked with † means the model is trained and evaluated by us.
Inference-free Sparse Retriever Sparse Retriever Dense Retriever
Dataset Our Model† BM25 SPLADE-doc-distill† SPLADE-v3-Doc SPLADE++SelfDistil SPLADE-v3-Distil ColBERTv2 Contriever TAS-B
TREC-COVID 72.4 68.8 68.4 68.1 71.0 70.0 73.8 59.6 48.1
NFCorpus 34.9 32.7 34.0 33.8 33.4 34.8 33.8 32.8 31.9
NQ 53.1 32.6 48.8 52.1 52.1 54.9 56.2 49.8 46.3
HotpotQA 67.9 60.2 62.6 66.9 68.4 67.8 66.7 63.8 58.4
FiQA-2018 36.4 25.4 31.2 33.6 33.6 33.9 35.6 32.9 30.0
ArguAna 49.1 47.2 37.7 46.7 47.9 48.4 46.3 44.6 42.9
Touche-2020 28.7 34.7 25.6 27.0 36.4 30.1 26.3 23.0 16.2
DBPedia-entity 40.5 28.7 35.9 36.1 43.5 42.6 44.6 41.3 38.4
SCIDOCS 16.7 16.5 14.7 15.2 15.8 14.8 15.4 16.5 14.9
FEVER 78.5 64.9 67.4 68.9 78.6 79.6 78.5 75.8 70.0
Climate-FEVER 19.2 18.6 15.1 15.9 23.5 22.8 17.6 23.7 22.8
SciFact 72.9 69.0 70.8 68.8 69.3 68.5 69.3 67.7 64.3
Quora 84.2 78.9 73.0 77.5 83.8 81.7 85.2 86.5 83.5
Average 50.35 44.48 45.02 46.97 50.56 49.99 49.95 47.54 43.67
The experiment results indicate that our model achieves SOTA in-domain performance among all inference-free retrievers. Additionally, the proposed model shrinks the performance gap between the inference-free retrievers and siamese sparse retrievers. Since there are trade-offs between search relevance and retrieval efficiency, we selected the Pareto optimal point based on the relationship curve between the two factors. It strikes a balance between search relevance and retrieval efficiency. Details can be found in Section 5.3.1.
5.2.2. Out-of-Domain Performance on BEIR
The purpose of out-of-domain(OOD) benchmarking is to test the model’s generalization capacity in a zero-shot fashion, which represents a quantity of real production scenarios, especially for ones with limited resource for building a relevance tuning pipeline. Following the work of Formal et al. (2022); Lassance et al. (2024), we evaluate our model on a readily available subset of 13 datasets from the BEIR benchmark, excluding CQADupstack, BioASQ, Signal-1M, TREC-NEWS, and Robust04. The comparison is shown in Table 4.
On the BEIR benchmark, our model’s zero-shot search relevance substantially outperforms other inference-free sparse retrievers, surpassing SPLADE-v3-Doc by a significant margin of 3.3 NDCG@10 score. Our model’s search relevance is comparable to SOTA siamese sparse retrievers and even outperforms strong siamese retrievers such as SPLADE-v3-DistilBERT and ColBERTv2. This result demonstrates that our model exhibits superior generalization capabilities. Moreover, we discovered that our model maintains stronger robustness in OOD settings compared to in-domain settings.
5.3. Retrieval Efficiency
5.3.1. Theoretical FLOPS number
FLOPS is the regularizer which controls the end-2-end efficiency of sparse retrieval. By adjusting $\lambda_{d}$ in equation 3, we can alter the degree of sparsity in the representation, as well as the computational cost incurred during the retrieval process  (Formal et al., 2021b; Paria et al., 2020). To investigate its impact, we conduct multiple sets of experiments by employing different values of $\lambda_{d}$ during fine-tuning, and measure the average FLOPS number for each query in BEIR datasets. The results are illustrated in Figure 2. The Pareto optimal point, which is regarded at the best trade-off between accuracy and efficiency, is marked in the diagram. In this paper, the corresponding checkpoint is used for other comparisons by default. Overall, in comparison to other inference-free sparse retrievers, our model exhibits not only superior search relevance but also enhanced retrieval efficiency. While siamese sparse retrievers possess search relevance comparable to or surpassing our model, our model’s retrieval efficiency is significantly superior, which presents a substantial advantage in production settings.
Figure 2. Search relevance vs efficiency on BEIR for sparse retrievers. Our models are trained with different $\lambda_{d}$.
5.3.2. End-to-end search performance
To evaluate the efficiency of our model in real production settings, the benchmark is conducted on a distributed OpenSearch cluster. The end-to-end search performance is measured, where end-to-end refers to the process of sending a raw text search request and marking it complete upon receiving the search response. All workloads are included, such as tokenizer inference, network traffic etc. We compare our method with BM25 in terms of the 99th percentile (P99) search latency and average search throughput under different concurrency levels by adjusting the client number. By default, BM25 in OpenSearch employs heuristic optimizations, such as block-max WAND (Ding and Suel, 2011), while learned sparse retrievers do not have these optimizations. We employ a simple heuristic optimization rule: first, we search for a preliminary result set using tokens with high IDF values, then we rerank the result set using all tokens [^11] The implmentation is based on OpenSearch rescore query.. The preliminary for this optimization rule is the involvement of IDF values. This optimization can boost search performance with negligible impact on search relevance. The results are listed in Table 5. Our method achieves an efficiency very close to that of BM25. The heuristic optimization rule boost the search latency and throughput for our method about 10 percentage. When both employ heuristic optimizations, the latency of our method is approximately 1.1x that of BM25.
Table 5. End-to-end search performance (milliseconds). Methods marked with † means the search process is optimized by heuristic rules.
Client-side P99 latency Mean throughput
Client # BM25† Ours Ours† BM25† Ours Ours†
5 13.4 21.7 17.6 784.2 484.8 586.2
10 20.9 25.2 22.9 1150.9 910.4 1024.5
20 35.4 38.2 38.7 1342.1 1183.4 1154.0
40 56.7 66.2 62.3 1658.6 1460.5 1537.7
80 74.7 91.7 81.1 2330.6 1858.19 2073.9
5.4. IDF-Aware FLOPS
5.4.1. Impact on search relevance
To assess the impact of IDF-aware FLOPS on the zero-shot search relevance, we conduct an ablation study employing different IDF settings on BEIR benchmark. In the default setting, we use fixed IDF values derived from MS MARCO dataset. We also conduct experiments using IDF values from the corresponding BEIR datasets. With different settings, we examined the impact of (1) employing IDF-aware FLOPS in training phase (2) retrieval with IDF values derived from different sources. We conduct these experiments on our model and SPLADE-doc-distill. The experiment results are shown in Table 6. From the experiment results, we obtain several conclusions: (1) The IDF-aware FLOPS boost the model search relevance at large margin. Training and inference with IDF, both our model and SPLADE-doc-distill achieve much better search relevance. (2) For models without pre-training phase, using IDF derived from the test set has better search relevance compared with fixed IDF derived from training data. However, if the model has undergone extensive pre-training on large-scale data using the fixed IDF, the conclusion is the opposite.
Table 6. The search relevance impact of IDF on BEIR dataset.
Model Trained Retrieval IDF avg NDCG
w/ IDF fixed BEIR
Ours 50.35
Ours 50.01
Ours w/o IDF $\times$ 48.65
SPLADE-doc-distill $\times$ 45.02
SPLADE-doc-distill + IDF 48.61
SPLADE-doc-distill + IDF 49.13
SPLADE-v3-doc $\times$ 46.97
5.4.2. Impact on retrieval efficiency and index size
Since the IDF-aware FLOPS applies larger penalty on low IDF tokens, experiments are conducted to measure its impact on retrieval efficiency. We conducted experiments to quantify the relationship between the expansion rate(average token number per expanded documents), retrieval efficiency(FLOPS number), and search relevance for our model. The comparison is made between our model with IDF-aware FLOPS and our model trained without IDF-aware FLOPS. The experimental results are depicted in Figure 3. The experimental findings demonstrate that the IDF-aware FLOPS substantially augments the retrieval efficiency. For models with similar expansion rate, the FLOPS number of those trained with IDF-aware FLOPS is much smaller.
Figure 3. The average FLOPS number, average token number per document and search relevance on BEIR datasets for models with and without IDF-aware FLOPS. The models are trained with different $\lambda_{d}$.
5.5. Heterogeneous Knowledge Distillation
We conducted an ablation study on the components of our proposed heterogeneous knowledge distillation to demonstrate their effectiveness. The detailed results are shown in Table 7, yielding several notable findings: Knowledge distillation is a more effective optimization objective than the naive InfoNCE loss. Utilizing supervision signals from one or more teacher models brings at least a 0.86 pts improvement (TAS-B) compared to model without pre-training, while the InfoNCE loss only improves 0.31 pts. Assembling multiple teachers consistently show further improvements, no matter for dense teacher models (+0.44 pts) or sparse ones (+0.18 pts). Our proposed heterogeneous knowledge distillation further enhances performance by 0.30 pts, amounting to a significant improvement of 1.74 pts. The process of normalizing outputs is crucial for model ensemble. The performance of using a simple additive ensemble method is close to that of a single teacher model (49.56). The significant performance drop empirically demonstrates the importance of addressing the scaling issue mentioned earlier.
Table 7. The search relevance impact of components in the pre-training stage.
Description BEIR avg score
Pre-training techniques
without pre-training 48.61
InfoNCE loss 48.92
ensemble KD, simply add 49.56
ensemble KD, norm and add (Ours) 50.35
Teacher models dense-only
TAS-B 49.47
gte-large 49.48
TAS-B & gte-large 49.92
Teacher models sparse-only
opensearch-sparse 49.66
SPLADE++ED 49.87
opensearch-sparse & SPLADE++ED 50.05
6. Conclusion
In this paper, we proposed two novel approaches to significantly improve the search relevance of inference-free learned sparse retrievers while maintaining high efficiency. We introduced IDF-aware FLOPS to mitigate underestimation of important tokens and increase sparsity, boosting both relevance and efficiency. We also developed a heterogeneous ensemble knowledge distillation framework leveraging strong dense and sparse retrievers for pre-training, enhancing generalization. Extensive experiments validate our methods’ effectiveness. Our model achieves SOTA performance among inference-free sparse retrievers on BEIR, while maintaining end-to-end latency only 1.1x that of BM25.